The daily production P in an automobile assembly plant is within 20 units of 500 units. Express the daily production as an absolute value inequality.
Answer by stanbon(75887)

The daily production P in an automobile assembly plant is within 20 units of 500 units. Express the daily production as an absolute value inequality.
----
Find the midpoint: (500+20)/2 = 260
Find the distance from midpoint to endpoint: 500-260 = 140
-----
Express that distance as an absolute expression:
|x - 260| <= 140
======================
Cheers,
Stan H.